# GQE: Generalized Query Expansion for Enhanced Text-Video Retrieval

### v1

@@ A.3 GQE Testing as a Training-free Tool
X-Pool [23] (Reproduced) 46.7 73.0 83.3 2.0 14.2
X-Pool [23] + GQE Testing 50.6 (+3.9) 76.3 86.5 1.0 10.4
CLIP-VIP [68] (Reproduced) 49.3 74.8 84.9 2.0 13.4
CLIP-VIP [68] + GQE Testing 53.2 (+3.9) 79.3 88.2 1.0 10.2
CLIP-VIP [68] + DSL [12] (Reproduced) 55.1 78.9 86.6 1.0 11.1
CLIP-VIP [68] + DSL [12] + GQE Testing 58.9 (+3.8) 83.4 89.5 1.0 8.357
@@
The text-to-video retrieval results are reported in Tab. 9. We observe that GQE Testing consistently improves the retrieval performance across the four methods. For the zero-shot CLIP, GQE Testing boosts the R@1 by 5.2% and R@5 by over 10%, which is a significant improvement. Notably, for the state-of-the-art CLIP-VIP [68], despite the remarkable results achieved, GQE Testing still improves the performance by a large margin, further validating the effectiveness of our method.
@@
We note that several popular methods, including DSL [12] and QB-Norm [7], enhance pre-trained text-video retrieval performance using a similar training-free approach. In the final experiment of Tab. 9, we demonstrate that our GQE Testing is compatible with these training-free methods. The combination of our approach and DSL significantly elevates the performance to a much higher level. Moreover, it is important to note that DSL [12] requires access to all queries during testing, which is an impractical assumption for real-world retrieval systems. In contrast, our approach does not require access to other queries during testing, making it more practical in real-world applications.
A.4 Video-to-Text Retrieval
In our work, we mainly focus on the task of text-to-video retrieval. In this section, we investigate the feasibility of applying generalized query expansion to the task of video-to-text retrieval. In this task, the query is a video while the retrieval targets are texts. Therefore, we apply the video captioning strategy described in our paper as query expansion. For each video query, we generate a set of captions to serve as auxiliary queries. We conduct retrieval for each query independently and aggregate the results via majority voting. The results are reported in Tab. 10. We observe that simply incorporating video captions as query expansion does not improve the performance of video-text retrieval.
In previous work [62], a common approach of processing captions is to sort and filter the captions according to the similarity scores between the video and its captions, since captions are often considered noisy. Inspired by this, we conduct experiments with various approaches for query (caption) selection. Nearest/farthest neighbor sampling sorts the captions based on the similarity between video and it captions, then selects the top-$k$ most similar/dissimilar captions as the expanded queries. FQS (Farthest Query Sampling) is described in our paper. We set $k$ to 2 in the experiments. The results demonstrate that the combination of video captioning as query expansion and the FQS algorithm effectively boosts the performance of video-text retrieval. In contrast, nearest/farthest neighbor sampling is shown to be not helpful. Upon comparing different query selection methods, we attribute the superiority of FQS to its ability to cover a larger and diverse area in the embedding space. The potential noise contained in the captions are effectively eliminated by majority voting.
A.5 Understanding Oracle Query Selection
The oracle query selection is implemented as follows: for one specific instance, among 1 GT query and $k$ expanded queries, we always select the query that yields the best ranking, where the best ranking refers to that the target video is ranked highest. If there are multiple such queries, we prioritize the GT query. We also provide the code in the supplementary material. Following the rule above, we stats the percentage of GT query vs. expanded query in the selected oracle query. The statistic is conducted on $1,000$ test samples of MSR-VTT dataset, $k=10$. The result shows that $59.5\%$ of selected queries come from the GT query set, while $40.5\%$ of selected queries are from the expanded queries. This actually provides an explanation of the huge performance gap between baseline and oracle, suggesting that the text queries can be specially investigated to enhance the retrieval.
Table 10: Video-to-text retrieval results on MSR-VTT [65].
Method R@1 R@5 R@10 MdR MnR
MMT [21] 27.0 57.5 69.7 3.7 21.3
T2VLAD [60] 31.8 60.0 71.1 3.0
SupportSet [42] 28.5 58.6 71.6 3.0 -
TMVM [33] 34.8 63.8 73.7 3.0 -
CLIP4Clip [38] 42.7 70.9 80.6 2.0 11.6
CenterCLIP [71] 42.8 71.7 82.2 2.0 10.9
CLIP2Video [17] 43.5 72.3 82.1 2.0 10.2
X-Pool (Baseline) [23] 44.4 73.3 84.0 2.0 9.0
TS2-Net [36] 45.3 74.1 83.7 2.0 9.2
DRL [58] 45.3 73.9 83.3 2.0 -
UATVR [16] 46.0 73.7 82.8 2.0 8.7
Cap4Video [62] 47.1 73.7 84.7 2.0 8.7
TEFAL [24] 47.1 75.1 84.9 2.0 7.4
Baseline 44.6 73.3 84.0 2.0 8.8
Baseline + GQE (All Captions) 42.0 64.6 73.6 2.0 55.0
Baseline + GQE (Nearest-Neighbor) 44.9 71.0 79.5 2.0 29.6
Baseline + GQE (Farthest-Neighbor) 45.9 71.9 80.7 2.0 24.4
Baseline + GQE (FQS) 52.2 78.3 86.5 1.0 7.9
A.6 Prompt for Testing Query Expansion
We provide the prompt used for generating testing queries in Fig. 4.
Figure 4: Prompts used for query generation.
A.7 Discussion, Limitation and Further Work
@@
Despite the significant performance achieved by our method, it exhibits certain limitations that serve as inspiration for future research. Firstly, in the query expansion of the training set, our approach focuses on a data-centric approach that expands the dataset. However, there is potential for further improvement by designing dedicated model architectures tailored to the expanded dataset, such as multi-instance learning. Secondly, in the query expansion of testing, we employ LLM as the tool to generate diverse queries. Given the vast prompting space of LLM, we have not invested significant efforts in prompt engineering. As a future research avenue, more dedicated prompts, such as in-context learning, can be harnessed to further enhance the quality of text queries. Finally, we acknowledge that there is still a noticeable gap in terms of retrieval performance and computational cost between our method and the upper bound of query expansion. This discrepancy encourages further research to bridge the gap.
A.8 Qualitative Examples
@@
In this section, we present a series of text-video retrieval examples from the MSR-VTT dataset [65]. These examples not only qualitatively validate the effectiveness of our approach but also demonstrate how the expanded queries aid in text-video retrieval. The examples are shown in Fig. 5, Fig. 7, Fig. 6, and Fig. 8. In each example, we show that the original query fails to retrieve the target video, while the expanded query does. Detailed explanations are presented in the caption of each figure, respectively.
@@
Figure 5: Rephrasing concepts. The original query focuses on the concept ‘team’. Although the concept of ‘team’ may imply the involvement of ‘multiple players’, it is not explicitly stated. The expanded query provides a more explicit view to elaborate the query, correcting the retrieval.
@@
Figure 6: Altering word order. The initially recalled video successfully covers the concept of ‘woman’, ‘man’, and ‘cooking food’. While it does not contain content corresponding to ‘setting a table’. Our hypothesis is that this is because ‘setting a table’ is positioned far back in the original query, limiting its impact. The expanded query recalls the video that captures the most terms in the query. This is achieved by altering the word order and emphasis of the original query.
@@
Figure 8: Robustness. In this example, the original query is clear while the initially recalled video appears to be not strongly relevant to the original query. This might be due to the inherent weakness of the retrieval model. However, the expanded query can still successfully retrieve the target video. This can be seen as an indication that query expansion improves the robustness of the retrieval model, making it resilient to corner failure cases.

### v2

@@ A.6 Text Enrichment in Retrieval as a Training-free Tool
X-Pool Gorti et al. (2022) (Reproduced) 46.7 73.0 83.3 2.0 14.2
X-Pool Gorti et al. (2022) + Text Enrich. Retrieval 50.6 (+3.9) 76.3 86.5 1.0 10.4
CLIP-VIP Xue et al. (2022) (Reproduced) 49.3 74.8 84.9 2.0 13.4
CLIP-VIP Xue et al. (2022) + Text Enrich. Retrieval 53.2 (+3.9) 79.3 88.2 1.0 10.2
CLIP-VIP Xue et al. (2022) + DSL Cheng et al. (2021) (Reproduced) 55.1 78.9 86.6 1.0 11.1
CLIP-VIP Xue et al. (2022) + DSL Cheng et al. (2021) + Text Enrich. Retrieval 58.9 (+3.8) 83.4 89.5 1.0 8.357
@@
The text-to-video retrieval results are reported in Tab. 12. We observe that Text Enrichment in Retrieval consistently improves the retrieval performance across the four methods. For the zero-shot CLIP, our method boosts the R@1 by 5.2% and R@5 by over 10%, which is a significant improvement. Notably, for the state-of-the-art CLIP-VIP Xue et al. (2022), despite the remarkable results achieved, our method still improves the performance by a large margin, further validating its effectiveness.
@@
We note that several popular methods, including DSL Cheng et al. (2021) and QB-Norm Bogolin et al. (2022), enhance pre-trained text-video retrieval performance using a similar training-free approach. In the final experiment of Tab. 12, we demonstrate that our Text Enrichment in Retrieval is compatible with these training-free methods. The combination of our approach and DSL significantly elevates the performance to a much higher level. Moreover, it is important to note that DSL Cheng et al. (2021) requires access to all queries during testing, which is an impractical assumption for real-world retrieval systems. In contrast, our approach does not require access to other queries during testing, making it more practical in real-world applications.
Figure 4: Prompts used for text enrichment in retrieval.
A.7 Prompt for Text Enrichment in Retrieval
We provide the prompt used for enriching testing retrieval queries in Fig. 4.
A.8 Limitation and Further Work
@@
Despite the significant performance achieved by our method, it exhibits certain limitations that serve as inspiration for future research. Firstly, in the text enrichment in training, our approach focuses on a data-centric approach that enriches the dataset. However, there is a potential for further improvement by designing dedicated model architectures tailored to the enriched dataset, such as multi-instance learning. Secondly, in the text enrichment of testing, we employ LLM as the tool to generate diverse queries. Given the vast prompting space of LLM, we have not invested significant efforts in prompt engineering. As a future research avenue, more dedicated prompts, such as in-context learning, can be harnessed to further enhance the quality of text queries. Thirdly, although our query selection mechanism has effectively reduced the computational cost in retrieval, the computation of LLM still exists. Optimization on this part can further benefit more practical applications. Fourthly, the involvement of VLMs and LLMs would inevitably introduce hallucinated content, which may limit the performance of the method. Although the mechanisms in the paper are shown to be resilient to noise, reducing hallucination has the potential to further improve the performance in the future. Finally, we acknowledge that there is still a noticeable gap in terms of retrieval performance and computational cost between our method and the upper bound. This discrepancy encourages further research to bridge the gap.
A.9 Qualitative Examples
@@
In this section, we present a series of text-video retrieval examples from the MSR-VTT dataset Xu et al. (2016b). These examples not only qualitatively validate the effectiveness of our approach but also demonstrate how the enriched queries aid in text-video retrieval. The examples from training set are shown in Fig. 9. The examples from the testing phase are shown in Fig. 5, Fig. 6, Fig. 7, and Fig. 8. In each example, we show that the original query fails to retrieve the target video, while the enriched query does. Detailed explanations are presented in the caption of each figure, respectively.
@@
Figure 5: Rephrasing concepts. The original query focuses on the concept ‘team’. Although the concept of ‘team’ may imply the involvement of ‘multiple players’, it is not explicitly stated. The enriched query provides a more explicit view to elaborate the query, correcting the retrieval.
@@
Figure 6: Altering word order. The initially recalled video successfully covers the concept of ‘woman’, ‘man’, and ‘cooking food’. While it does not contain content corresponding to ‘setting a table’. Our hypothesis is that this is because ‘setting a table’ is positioned far back in the original query, limiting its impact. The enriched query recalls the video that captures the most terms in the query. This is achieved by altering the word order and emphasis of the original query.
@@
Figure 8: Robustness. In this example, the original query is clear while the initially recalled video appears to be not strongly relevant to the original query. This might be due to the inherent weakness of the retrieval model. However, the enriched query can still successfully retrieve the target video. This can be seen as an indication that text enrichment improves the robustness of the retrieval model, making it resilient to corner failure cases.
Figure 9: Examples from MSR-VTT training dataset with text enrichment.